OpenStudy (anonymous):

Simplify. 10+3x-x^2/x^2+6x-7=

OpenStudy (anonymous):

first lets order the terms \[\frac{ -x^2+3x+10 }{ x^2+6x-7 }\] so do u know how to factor

OpenStudy (anonymous):

yes I do.

OpenStudy (anonymous):

so what does the top factor out into

OpenStudy (anonymous):

would you have to factor out a -1 first so that the numerator has a positive a value?

OpenStudy (anonymous):

yes i would do that first :)

OpenStudy (anonymous):

I got \[\frac{ -1(x-5)(x+2) }{ (x+1)(x-7) }\]

OpenStudy (anonymous):

oh oops that's wrong.

OpenStudy (anonymous):

should have +7 and - 1

OpenStudy (anonymous):

but other than that - you got it

OpenStudy (anonymous):

yeah that's what I thought haha.. but would that be the final answer? In the answer sheet it says the right answer would be \[\frac{ x+2 }{ x+1 }\]

OpenStudy (anonymous):

did u copy the question down right look back at the original ?

OpenStudy (anonymous):

Yeah I did.. hmmm